OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

jigglypuff314 (jigglypuff314):

\[slope = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2) standard form = ax + by = c first you would have to get the equation of the line in either slope-intercept form or point-slope form, which were you taught?

OpenStudy (anonymous):

both but I still dont understand

jigglypuff314 (jigglypuff314):

can you get the slope?

OpenStudy (anonymous):

could one of you guys help with my question?

OpenStudy (anonymous):

please?

jigglypuff314 (jigglypuff314):

i don't see why not

OpenStudy (anonymous):

thx! its about matrices :P

jigglypuff314 (jigglypuff314):

umm... then Idk sorry :P

OpenStudy (anonymous):

haha! your fine

OpenStudy (anonymous):

jigglypuff314, is it -2/-3

jigglypuff314 (jigglypuff314):

@Gregk but it's just logic... what numbers in the table in the august column match the matrix?

jigglypuff314 (jigglypuff314):

yes and negative of negative is positive so the slope is 2/3

jigglypuff314 (jigglypuff314):

then point-slope form is: y-y1 = m(x-x1) and you plug in what you know into that

OpenStudy (anonymous):

@jigglypuff314 feeling stupid lol

OpenStudy (anonymous):

thanks. you are a life saver :D

jigglypuff314 (jigglypuff314):

so you know how to do it now @Ummah.Naja ?

OpenStudy (anonymous):

Yes Sir. Thanks

jigglypuff314 (jigglypuff314):

okay, your welcome, and feel free to ask me if you have any more questions :)

OpenStudy (anonymous):

Okey Dokey, Artichokey

OpenStudy (anonymous):

I apologize for that lame, corny reply. lol lml : )